# Improving Ground State Accuracy of Variational Quantum Eigensolvers with Soft-coded Orthogonal Subspace Representations

### v1

@@ I Introduction
Assumption A.1 is particularly relevant for low-symmetry systems such as general realizations of spin glasses, which are usually mentioned as potential candidates for quantum advantage, especially in the context of quantum annealing approaches PhysRevX.5.031026; tanaka2017quantum; chowdhury2025pushing. These type of systems motivates the analysis of the approach without using any information about possible symmetries. Therefore, for testing purposes, we decided to ignore information even in the case of the Transverse-Field Ising (TFI) model in 2D, which we consider as benchmark case in Section III.2. Another motivation for considering non-symmetric ansätze is that enforcing problem-specific symmetries into the ansatz makes it typically deeper and possibly not feasible for NISQ era applications. Assumption A.2, on the other hand, makes the analysis and comparison between different subspace representations fairer. This does not prevent future extensions and integrations of the approach proposed here with other approaches based on sequential improvements of the ansatz (relaxing assumption A.2) or with the exploitation of system-specific properties (relaxing assumption A.1), but these possible variations would not enrich the discussion on the advantages of the soft-coded representation and are left for future investigations.
@@ II.4 Subspace Expressibility
where $V$ is a (unitary) linear combination of generally non-commuting unitaries. This representation can therefore generate highly entangled states even with shallow circuits for $U_{p}(\bm{\theta}_{p})$, enriching the set of reachable states and therefore the expected expressibility. However, while this intuition seems reasonable and is supported by numerical evidence, as shown in Section III, a quantitative analysis of the expressibility of such representations is not straightforward, since it would require defining proper metrics for the expressibility of single states within a parametric subspace, which we did not find in literature. Indeed, most of the literature on expressibility and its relation with barren plateaus focuses on single state ansätze ExpressibilitySim2019; Funcke2021dimensional; Arrasmith_2022; PRXQuantum.3.010313; Brozzi:2025omu; often depending on comparison with Haar-random states or t-designs, which are not directly applicable to subspace representations, since a non-uniform coverage of the full Hilbert space is not an issue, as long as the relevant subspace (e.g., low-energy states) is well represented by at least some portion of the parameterized subspaces $\cal V_{\bm{\theta}}$. As possible metric, we considered a random sampling of subspaces $\cal V_{\bm{\theta}}$ for random parameters $\bm{\theta}$, which would provide an approximation to the relative coverage of the Hilbert space $\cup_{\bm{\theta}}{\cal V}_{\bm{\theta}}$ but, in practice, this approach is unfeasible even for systems with very few qubits. For these reasons, a quantitative analysis of the expressibility of the considered subspace representations is left for future work.
@@ III.1 Numerical Setup
Regarding the optimization loop, we used the NFT optimizer Nakanishi:2019rrm from Qiskit Algorithms, but different choices would mostly affect the convergence rates, not the final fidelities achieved. The parameters for different runs are initialized independently and randomly from a uniform distribution in the reduced-domain $0.2\pi,0.2\pi]$. This choice aligns with common practices in VQE literature, aiming to avoid potential issues such as barren plateaus by restricting the initial parameters to a smaller range, as well as ensuring an appreciable overlap with the target state, which is shown to be quite relevant for the convergence of VQE-like algorithms zhang2022escaping; PhysRevApplied.22.054005; Arrasmith_2022; PRXQuantum.3.010313; Larocca:2024plh. In all cases considered, we set the maximum number of iterations to $1500$. Depending on the total number of parameters, this might not always be sufficient to reach convergence, as we show in the following. Furthermore, at each step of the optimization we also compute fidelity metrics (both subspace fidelity and $H_{\rm trc}$ fidelity, as defined in Section II.3), which involve using the exact ground state and performing exact diagonalization of the Hamiltonian restricted to the subspace. We stress that these intermediate measurements are not used in the optimization process itself, but are useful for monitoring more fine-grained aspects of the convergence behavior besides the global loss. In particular, for the subspace representations, the matrix elements $\langle{\psi_{p}}|{H}|{\psi_{q}}\rangle$} (and the overlap matrix elements $\left\langle\psi_{p}\middle|\psi_{q}\right\rangle$, in the soft-ortho case) are meant to be measured only as final stage of the run after convergence is reached.
@@ III.2.1 Relation Between Cost and Fidelity During Optimization
As mentioned for Figure 4, some runs of standard VQE (and hard-ortho case, in lesser extent) exhibit a non-monotonic behavior where the cost function decrease is accompanied by an increase in infidelity. This indicates that the cost function alone is not a good surrogate for fidelity in these cases, especially in the latest stage of the optimization. On the other hand, soft-ortho runs seem to follow more direct paths towards higher fidelities. It is also interesting to notice that, while the initial infidelity values of both the standard VQE and the hard-ortho runs are relatively large (i.e., low overlap with the ground state), most of the the runs in the soft-ortho representation start with already reasonably low infidelities. Regarding the pairwise overlap $\left|\left\langle\psi_{q}(\bm{\theta}_{(q)})\middle|\psi_{p}(\bm{\theta}_{(p)})\right\rangle\right|^{2}$ between distinct frame states from the soft-ortho representation, we monitored their value during the optimization. At the beginning of the run, the random initialization of the parameters does not make it meaningful, while at the end of the optimization, where the penalty term takes effect over the relative energy gaps, we never measured values exceeding $15\%$ for all subspace dimensions $K$ and number of layers $N_{l}$ investigated. Also, it can reach values up to $30\%$ during the run (besides the very first few iterations). Moreover, for $K=2$ the overlap between the two frame states, while possibly peaking during the optimization up to $20\%$, reaches overlaps lower than $5\%$ at the end of the run.
@@
K Layers ($N_{l}$) Max (%) Med (%) Max (%) Med (%)
2 2 - - 0.003 0.220
4 0.010 0.181 0.010 0.050
6 0.395 0.661 0.002 0.135
8 0.318 0.815 0.034 0.050
10 0.046 0.846 - -
12 0.017 0.011 - -
4 2 - - 0.150 0.232
4 0.655 1.034 0.011 0.075
6 1.025 0.918 0.041 0.254
8 0.786 0.858 0.068 0.201
10 0.950 0.991 - -
12 0.433 0.980 - -
8 2 - - 0.089 0.076
4 1.156 1.617 0.119 0.082
6 1.279 1.923 0.254 0.404
8 1.635 1.705 0.202 0.221
Table 1: Discrepancy $\Delta{\cal F}={\cal F}_{\rm sub}-{\cal F}_{\rm trc}$ between the subspace (Eq. (12)) and $H_{\rm trc}$ fidelity (Eq. (13)) definitions for either best (Max) or median (Med) values across 10 independent runs for each fixed number layers ($N_{l}$) or subspace dimension $K$ for the transverse-field Ising model on a $3\times 3$ lattice. Missing values are flagged with a hyphen and correspond to simulation not performed.
Since the fidelity data reported in Figures 4 and 5 refer only to the $H_{\rm trc}$ definition, i.e., the fidelity of the state obtained from the diagonalization of the truncated Hamiltonian, one might wonder how this compares with the subspace fidelity $\mathcal{F}_{\rm sub}$, which measures the overlap of the exact ground state with the subspace itself, as introduced in Section II.3. To this end, in Table 1 we report the discrepancy $\Delta{\cal F}:=\mathcal{F}_{\rm sub}-\mathcal{F}_{\rm trc}$ at the end of the optimization runs for both hard-ortho and soft-ortho representations, computing it between either the best or the median fidelities across the 10 independent runs for each ansatz configuration. In all cases considered, the discrepancy never exceeds $2\%$ for the hard-ortho and $0.5\%$ for the soft-ortho representation. This observation confirms that, at least for the present model and ansatz structures, the subspace at convergence ${\cal V}_{\bm{\theta}^{*}}$ is well aligned with an eigensector of the Hamiltonian (possibly the low-energy one), with relatively negligible mixing with the complementary subspace $\overline{\cal V}_{\bm{\theta}^{*}}$. Therefore, the best estimate obtained from the diagonalization of the truncated Hamiltonian is not too far (within the percent level) from the best possible state that can be obtained from the variational subspace itself.
@@
The behavior of the cumulative infidelity for the three approaches considered and at a common fixed number of 144 parameters is shown in Figure 6 for randomly picked individual runs [^5] Threshold energies including the ground state $\mu\leq E_{0}$ are excluded from Figure 6, since ${\mathbb{P}}_{\mu\geq E_{0}}=\mathds{1}$ and they trivially correspond to ${\rm CI}=0$.. At first, all state representations have a significant overlap with high-energy eigenstates, while they progressively reduce it in favor of low-energy contributions as the optimization proceeds. The standard VQE and hard-ortho representation show a similar behavior, with significant overlaps with mid-energy eigenspaces even at the end of the optimization. It is also worth looking, especially for the standard VQE panel, the very bottom side strip of the spectrum involving the first excited eigenspace above the ground state. Here we see that, as the overlap with higher-energy states continues to deplete as the optimization run progresses, the overlap including the first excited eigenspace shows a non-monotonic behavior: up to intermediate iterations (around $\sim 600$) it lowers but, going further in the optimization (i.e., iterations $\gtrsim 1000$) the contamination from the first excited eigenspace increases again. This indicates that, while optimizing the average energy of the variational state, the overall overlap with high-energy states is reduced but, depending on the expressivity of the ansatz and accessible regions of the parameter space, the overlap with the above-ground eigenspaces is not prevented from increasing, especially at convergence where energy variations are smaller and the dynamics of the optimization is more constrained by the ansatz structure. This is in line with the non-monotonic behavior of the fidelity observed in Figure 4 for standard VQE. On the other hand, the soft-ortho representation shows a more effective suppression not only of high-energy contributions, but also of mid-energy and just-above-ground contributions, leading to a more robust convergence to the ground state. Indeed, realizing that this very non-monotonic behavior of the ground state fidelity is at play clarifies also why approaches involving multiple states might be more effective even in tasks targeting only the ground state. The hard-ortho approach however, while showing some improvements over standard VQE, and mostly preventing this behavior, is still not as effective as the soft-ortho one in reaching high fidelities, possibly due to the more stringent constraints on the expressivity of the variational frame, as argued in Section II.4.
@@ III.2.2 Fidelity Dependence on Subspace Dimension and Number of Parameters
A summary of results for the final best and median $H_{\rm trc}$ fidelities (after 1500 iterations) for the three approaches considered and different ansatz depths and subspace dimension $K$ is reported in Table 2, where the general behavior discussed above is confirmed. The cases with larger subspace dimension $K$ seems to be more challenging for both hard-ortho and soft-ortho representations, with a slight decrease in the best fidelities achieved at fixed number of layers, which seems to indicate a generally slower convergence. To get a better picture of the general dependence of the final fidelity, we provide two point of views: one relating the final fidelity ${\cal F}_{\rm trc}$ with the subspace dimension $K$ (Figure 7), and another one relating it with the number of trainable parameters in the (subspace) ansatz (Figure 8). Both perspectives reflect the distribution of final fidelities over 10 independent runs, which are aggregated in the aforementioned Table 2, reporting only the maximum and the median fidelities.
@@
Let us discuss first Figure 7, where we show the dependence of the distribution of final $\mathcal{F}_{\rm trc}$ fidelities over 10 independent runs and as a function of the subspace dimension $K$. Looking at the best fidelities achieved, we can see that both VQE and hard-ortho remain in the $75$-$90\%$ range, while soft-ortho reaches fidelities above $95\%$ for all values of $K$ and all numbers of layers in the ansätze. In particular, hard-ortho shows no improvement over standard VQE if looking at best runs, while a marginal improvement ($\sim 70\%$ to $\sim 80\%$) is observed when looking at median fidelities. For soft-ortho, we also see that increasing the subspace dimension beyond $K=2$ does not seem to improve the fidelity, or at least a possible marginal improvement is traded for a slower convergence which makes the cut-off of 1500 iterations insufficient. This is also evident in the hard-ortho case, even if its number of parameters at fixed number of layers is the same as standard VQE. From this observation, one might conclude that the advantage of a subspace approach, at least for the present model and ansatz structures, is mostly granted by the inclusion of just one additional state in the frame, while further increasing the subspace dimension does not provide significant benefits. This might be related to the fact that the first excited state is the one mostly contaminating the variational ground state estimate at the end of the optimization, as argued in Section III.2.1 for Figure 6; just including a second state to the variational frame seems to be already sufficient to make the optimal subspace ${\cal V}_{\bm{\theta}^{*}}$ significantly overlap with the exact first excited state, thus allowing in the diagonalization stage to discriminate ground state components with better accuracy. The second perspective we provide on the final fidelities achieved is reported in Figure 8 with an emphasis on the fixed number of trainable parameters. Soft-ortho clearly outperforms standard VQE and hard-ortho for all numbers of parameters explored and all runs. While standard VQE and hard-ortho approaches show a relatively large spread in the final fidelities across runs even, the final estimates obtained with soft-ortho are more robustly clustered around high-fidelity values, with quite smaller interquartile ranges, confirming a more reliable convergence behavior. Moreover, the worst run for the soft-ortho representation with only 72 parameters (i.e., $N_{l}=2$ layers and $K=2$) already exceeds the best fidelity obtained in both VQE and hard-ortho with all numbers of parameters considered. This confirms that shallower circuits are already sufficient to reach high accuracies when using the soft-ortho subspace representation. Indeed, as observed for Figure 7, deeper circuits with larger number of parameters seem to improve the median final fidelity just marginally and up to a point (around 144 parameters or $N_{l}=8$ layers), after which it becomes counterproductive, likely due to the cut-off imposed of maximum 1500 total iterations (as evident from the cost histories in Figure 3). In general, as argued in Section II.4, we observe that sacrificing some entanglement expressivity is not expected to quicken convergence but can be beneficial on reaching higher accuracies: frames with shallower circuits would reach lower levels of global entanglement, but considering general linear combinations of independent states in the frame allows to recover higher accuracies on the ground state estimates.

### v2

@@ I Introduction
Assumption A.1 is particularly relevant for low-symmetry systems such as general realizations of spin glasses, which are usually mentioned as potential candidates for quantum advantage, especially in the context of quantum annealing approaches PhysRevX.5.031026; tanaka2017quantum; chowdhury2025pushing. These types of systems motivate the analysis of the approach without using any information about possible symmetries. Therefore, in Section III.2, we consider as a benchmark the 2D Transverse-Field Ising (TFI) model, but without exploiting information on the system. Another motivation for considering non-symmetric ansätze is that enforcing problem-specific symmetries into the ansatz makes it typically deeper and possibly not feasible for NISQ era applications. Assumption A.2, on the other hand, makes the analysis and comparison between different subspace representations fairer. This does not prevent future extensions and integrations of the approach proposed here with other approaches based on sequential improvements of the ansatz (relaxing assumption A.2) or with the exploitation of system-specific properties (relaxing assumption A.1), but these possible variations would not enrich the discussion on the advantages of the soft-coded representation and are left for future investigations.
@@ II.4 Subspace Expressibility
where $V$ is a linear combination of generally non-commuting unitaries. This representation can therefore generate highly entangled states even with shallow circuits for $U_{p}(\bm{\theta}_{p})$, enriching the set of reachable states and therefore the expected expressibility. However, while this intuition seems reasonable and is supported by numerical evidence, as shown in Section III, a quantitative analysis of the expressibility of such representations is not straightforward, since it would require defining proper metrics for the expressibility of single states within a parametric subspace, which we did not find in literature. Indeed, most of the literature on expressibility and its relation with barren plateaus focuses on single state ansätze ExpressibilitySim2019; Funcke2021dimensional; Arrasmith_2022; PRXQuantum.3.010313; Brozzi:2025omu; often depending on comparison with Haar-random states or t-designs, which are not directly applicable to subspace representations, since a non-uniform coverage of the full Hilbert space is not an issue, as long as the relevant subspace (e.g., low-energy states) is well represented by at least some portion of the parameterized subspaces $\cal V_{\bm{\theta}}$. As possible metric, we considered a random sampling of subspaces $\cal V_{\bm{\theta}}$ for random parameters $\bm{\theta}$, which would provide an approximation to the relative coverage of the Hilbert space $\cup_{\bm{\theta}}{\cal V}_{\bm{\theta}}$ but, in practice, this approach is unfeasible even for systems with very few qubits. For these reasons, a quantitative analysis of the expressibility of the considered subspace representations is left for future work.
@@ III.1 Numerical Setup
Regarding the optimization loop, we used the NFT optimizer Nakanishi:2019rrm from Qiskit Algorithms, but different choices would mostly affect the convergence rates, not the final fidelities achieved. The parameters for different runs are initialized independently and randomly from a uniform distribution in the reduced-domain $0.2\pi,0.2\pi]$. This choice aligns with common practices in VQE literature, aiming to avoid potential issues such as barren plateaus, as well as ensuring an appreciable overlap with the target state, which is shown to be quite relevant for the convergence of VQE-like algorithms zhang2022escaping; PhysRevApplied.22.054005; Arrasmith_2022; PRXQuantum.3.010313; Larocca:2024plh. In all cases considered, we set the maximum number of iterations to $1500$. Depending on the total number of parameters, this might not always be sufficient to reach convergence, as we show in the following. Furthermore, at each step of the optimization we also compute fidelity metrics (both subspace fidelity and $H_{\rm trc}$ fidelity, as defined in Section II.3), which involve using the exact ground state and performing exact diagonalization of the Hamiltonian restricted to the subspace. We stress that these intermediate measurements are not used in the optimization process itself, but are useful for monitoring more fine-grained aspects of the convergence behavior besides the global loss. In particular, for the subspace representations, the matrix elements $\langle{\psi_{p}}|{H}|{\psi_{q}}\rangle$} (and the overlap matrix elements $\left\langle\psi_{p}\middle|\psi_{q}\right\rangle$, in the soft-ortho case) are meant to be measured only as final stage of the run after convergence is reached.
@@ III.2.1 Relation Between Cost and Fidelity During Optimization
As mentioned for Figure 4, some runs of standard VQE (and, to a lesser extent, the hard-ortho case) exhibit a non-monotonic behavior where the cost function decrease is accompanied by an increase in infidelity. This indicates that the cost function alone is not a good surrogate for fidelity in these cases, especially in the latest stage of the optimization. On the other hand, soft-ortho runs seem to follow more direct paths towards higher fidelities. It is also interesting to notice that, while the initial infidelity values of both the standard VQE and the hard-ortho runs are relatively large (i.e., low overlap with the ground state), most of the the runs in the soft-ortho representation start with already reasonably low infidelities. Regarding the pairwise overlap $\left|\left\langle\psi_{q}(\bm{\theta}_{(q)})\middle|\psi_{p}(\bm{\theta}_{(p)})\right\rangle\right|^{2}$ between distinct frame states from the soft-ortho representation, we monitored their value during the optimization. At the beginning of the run, the random initialization of the parameters does not make it meaningful, while at the end of the optimization, where the penalty term takes effect over the relative energy gaps, we never measured values exceeding $15\%$ for all subspace dimensions $K$ and number of layers $N_{l}$ investigated. Also, it can reach values up to $30\%$ during the run (besides the very first few iterations). Moreover, for $K=2$ the overlap between the two frame states, while possibly peaking during the optimization up to $20\%$, reaches overlaps lower than $5\%$ at the end of the run.
@@
$K$ Layers ($N_{l}$) $\max{\Delta{\cal F}}$ % ${\rm med}{\Delta{\cal F}}$ % $\max{\Delta{\cal F}}$ % ${\rm med}{\Delta{\cal F}}$ %
2 2 - - 0.3 <0.1
4 0.7 <0.2 0.1 0.02(1)
6 1.0 0.5(2) 0.6 <0.02
8 1.3 0.5(1) 0.2 <0.04
10 1.1 0.8(2) - -
12 1.1 <0.3 - -
4 2 - - 0.2 0.15(2)
4 0.9 0.73(5) 0.3 0.10(2)
6 1.0 0.76(4) 0.4 0.07(5)
8 0.8 0.70(5) 0.5 0.13(5)
10 1.2 0.78(8) - -
12 0.9 0.74(6) - -
8 2 - - 0.3 0.11(5)
4 1.3 1.25(1) 0.1 0.07(1)
6 1.6 1.21(5) 0.8 0.2(1)
8 1.4 1.29(3) 0.9 0.29(7)
[…]
Since the fidelity data reported in Figures 4 and 5 refer only to the $H_{\rm trc}$ definition, i.e., the fidelity of the state obtained from the diagonalization of the truncated Hamiltonian, one might wonder how this compares with the subspace fidelity $\mathcal{F}_{\rm sub}$, which measures the overlap of the exact ground state with the subspace itself, as introduced in Section II.3. To this end, in Table 1 we report [^5] Errors in tables are estimated with bootstrap resampling for quantities involving medians, while quantities involving extremal values (e.g., min, max) are reported without error (see Ref. Bickel1981; Angus01011992 for a discussion). the median and maximal discrepancy $\Delta{\cal F}:=\mathcal{F}_{\rm sub}-\mathcal{F}_{\rm trc}$ at the end of the optimization runs for both hard-ortho and soft-ortho representations, computing it across the 10 independent runs for each ansatz configuration. In all cases considered, the discrepancy never exceeds $2\%$ for the hard-ortho and $1\%$ for the soft-ortho representation. This observation confirms that, at least for the present model and ansatz structures, the subspace at convergence ${\cal V}_{\bm{\theta}^{*}}$ is well aligned with an eigensector of the Hamiltonian (possibly the low-energy one), with relatively negligible mixing with the complementary subspace $\overline{\cal V}_{\bm{\theta}^{*}}$. Therefore, the best estimate obtained from the diagonalization of the truncated Hamiltonian is not too far (within the percent level) from the best possible state that can be obtained from the variational subspace itself.
@@
The behavior of the cumulative infidelity for the three approaches considered and at a common fixed number of 144 parameters is shown in Figure 6 for randomly picked individual runs [^6] Threshold energies including the ground state $\mu\leq E_{0}$ are excluded from Figure 6, since ${\mathbb{P}}_{\mu\geq E_{0}}=\mathds{1}$ and they trivially correspond to ${\rm CI}=0$.. At first, all state representations have a significant overlap with high-energy eigenstates, while they progressively reduce it in favor of low-energy contributions as the optimization proceeds. The standard VQE and hard-ortho representation show a similar behavior, with significant overlaps with mid-energy eigenspaces even at the end of the optimization. It is also worth looking, especially for the standard VQE panel, the very bottom side strip of the spectrum involving the first excited eigenspace above the ground state. Here we see that, as the overlap with higher-energy states continues to deplete as the optimization run progresses, the overlap including the first excited eigenspace shows a non-monotonic behavior: up to intermediate iterations (around $\sim 600$) it lowers but, going further in the optimization (i.e., iterations $\gtrsim 1000$) the contamination from the first excited eigenspace increases again. This indicates that, while optimizing the average energy of the variational state, the overall overlap with high-energy states is reduced but, depending on the expressivity of the ansatz and accessible regions of the parameter space, the overlap with the above-ground eigenspaces is not prevented from increasing, especially at convergence where energy variations are smaller and the dynamics of the optimization is more constrained by the ansatz structure. This is in line with the non-monotonic behavior of the fidelity observed in Figure 4 for standard VQE. On the other hand, the soft-ortho representation shows a more effective suppression not only of high-energy contributions, but also of mid-energy and low-lying excited states contributions, leading to a more robust convergence to the ground state. Indeed, realizing that this very non-monotonic behavior of the ground state fidelity is at play clarifies also why approaches involving multiple states might be more effective even in tasks targeting only the ground state. The hard-ortho approach however, while showing some improvements over standard VQE, and mostly preventing this behavior, is still not as effective as the soft-ortho one in reaching high fidelities, possibly due to the more stringent constraints on the expressivity of the variational frame, as argued in Section II.4.
@@ III.2.2 Fidelity Dependence on Subspace Dimension and Number of Parameters
A summary of results for the final best and median $H_{\rm trc}$ fidelities (after 1500 iterations) for the three approaches considered and different ansatz depths and subspace dimension $K$ is reported in Table 2, where the general behavior discussed above is confirmed. The cases with larger subspace dimension $K$ seem to be more challenging for both hard-ortho and soft-ortho representations, with a slight decrease in the best fidelities achieved at fixed number of layers, which seems to indicate a generally slower convergence. To get a better picture of the general dependence of the final fidelity, we provide two points of views: one relating the final fidelity ${\cal F}_{\rm trc}$ with the subspace dimension $K$ (Figure 7), and another one relating it with the number of trainable parameters in the (subspace) ansatz (Figure 8). Both perspectives reflect the distribution of final fidelities over 10 independent runs, which are aggregated in the aforementioned Table 2, reporting only the maximum and the median fidelities.
@@
Let us discuss first Figure 7, where we show the dependence of the distribution of final $\mathcal{F}_{\rm trc}$ fidelities over 10 independent runs and as a function of the subspace dimension $K$. Looking at the best fidelities achieved, we can see that both VQE and hard-ortho remain in the $75$-$90\%$ range, while soft-ortho reaches fidelities above $95\%$ for all values of $K$ and all numbers of layers in the ansätze. In particular, hard-ortho shows no improvement over standard VQE if looking at best runs, while a marginal improvement ($\sim 70\%$ to $\sim 80\%$) is observed when looking at median fidelities. For soft-ortho, we also see that increasing the subspace dimension beyond $K=2$ does not seem to improve the fidelity, or at least a possible marginal improvement is traded for a slower convergence which makes the cut-off of 1500 iterations insufficient. This is also evident in the hard-ortho case, even if its number of parameters at fixed number of layers is the same as standard VQE. From this observation, one might conclude that the advantage of a subspace approach, at least for the present model and ansatz structures, is mostly granted by the inclusion of just one additional state in the frame, while further increasing the subspace dimension does not provide significant benefits. This might be related to the fact that the first excited state is the one mostly contaminating the variational ground state estimate at the end of the optimization, as argued in Section III.2.1 for Figure 6; including a second state to the variational frame seems to be already sufficient to make the optimal subspace ${\cal V}_{\bm{\theta}^{*}}$ significantly overlap with the exact first excited state, thus allowing in the diagonalization stage to discriminate ground state components with better accuracy. The second perspective we provide on the final fidelities achieved is reported in Figure 8 with an emphasis on the fixed number of trainable parameters. Soft-ortho clearly outperforms standard VQE and hard-ortho for all numbers of parameters explored and all runs. While standard VQE and hard-ortho approaches show a relatively large spread in the final fidelities across runs, the final estimates obtained with soft-ortho are more robustly clustered around high-fidelity values, with significantly smaller interquartile ranges, confirming a more reliable convergence behavior. Moreover, the worst run for the soft-ortho representation with only 72 parameters (i.e., $N_{l}=2$ layers and $K=2$) already exceeds the best fidelity obtained in both VQE and hard-ortho with all numbers of parameters considered. This confirms that shallower circuits are already sufficient to reach high accuracies when using the soft-ortho subspace representation. Indeed, as observed for Figure 7, deeper circuits with larger number of parameters seem to improve the median final fidelity just marginally and up to a point (around 144 parameters or $N_{l}=8$ layers), after which it becomes counterproductive, likely due to the cut-off imposed of maximum 1500 total iterations (as evident from the cost histories in Figure 3). In general, as argued in Section II.4, we observe that sacrificing some entanglement expressivity is not expected to quicken convergence but can be beneficial on reaching higher accuracies: frames with shallower circuits would reach lower levels of global entanglement, but considering general linear combinations of independent states in the frame allows to recover higher accuracies on the ground state estimates.